# Interacting Multiple Model-based Joint Homography Matrix and Multiple Object State Estimation

### v1

@@ 5 Experiments
For all experiments and test set results, the camera calibration matrices provided in [14] are used – except for the KITTI dataset, which provides its own. Furthermore, the estimates of $\mathbf{A}_{t}$ (2) provided by [14] are used for the MOT17 and MOT20 datasets, but for the DanceTrack and KITTI datasets optical flow features and the RANSAC algorithm are used to estimate $\mathbf{A}_{t}$ with the OpenCV library.
@@
For IMM-JHSE, the following hyperparameters need to be specified: $\sigma_{x},\sigma_{y},\alpha_{1},\alpha_{2},\alpha_{3},\Omega,b,d_{\text{high}},d_{\text{low}}$, as well as the transition probabilities illustrated in Figure 5. For the transition probabilities, it is only necessary to specify the self-transition probabilities $p_{s,s},p_{d,d},p_{G,G}$ and $p_{I,I}$, since the other transition probabilities can be derived from them.
@@
This paper uses the pattern search algorithm [54] to find optimum values for the hyperparameters. For each experiment described subsequently, the initial values are set as follows: $\sigma_{x}=5,\sigma_{y}=5,\alpha_{1}=\alpha_{2}=\alpha_{3}=0.5,\Omega=30,b=0,d_{\text{high}}=0.6,d_{\text{low}}=0.5,p_{s,s}=p_{d,d}=p_{G,G}=p_{I,I}=0.9$, and the number of pattern search iterations is restricted to 200. The objective function selected for optimisation is HOTA for the MOT17, MOT20 and KITTI datasets and association accuracy (AssA) for the DanceTrack dataset. Other evaluation metrics include IDF1 and multiple object tracking accuracy (MOTA). The reader is referred to [55] for details regarding these metrics. In all cases, evaluation is performed with the TrackEval library [56]. For the DanceTrack, KITTI and MOT20 datasets, the combined objective function of the entire dataset is optimised, while for the MOT17 dataset, the objective function is optimised per video.
@@
Table 1 reports the results of an ablation study on the DanceTrack validation dataset. The second column refers to the use of the dynamic measurement noise estimation method described in Section 4.3. When dynamic measurement noise is not used (as indicated with X), (14) is used to provide an estimate of the measurement noise at each time step. When the dynamic estimation method is not used, HOTA decreases from $63.81$ to $60.80$. Thus, the dynamic noise estimation method contributes significantly to the success of IMM-JHSE. The third column refers to the use of an IMM filter. Without an IMM filter, IMM-JHSE is restricted to the single camera motion model of (8). In this case, HOTA decreases to $62.04$ – showing that the mixture with the static camera motion model (9) allows for more accurate state propagation. The final column refers to the use of the image plane filter and the BIoU. Without these, all association scores are calculated by $P(D)d_{\text{conf}}$, and HOTA decreases to $63.03$. Thus, the use of bounding box information in conjunction with the ground plane state vector proves beneficial.
@@
In Table 2, the way bounding box information is considered is varied. For “IMM-like”, the association score for stages 2 and 3 remains as in (21). For “Multiplicative”, all stages use the association score in (20). The “IMM-like” association scores achieve the best performance across all metrics for both MOT17 and DanceTrack. Furthermore, the benefit of using the “IMM-like” association score is illustrated in Figure 7, where track identities are maintained while the performers jump, and the ground plane position estimates cannot be relied upon. Thus, the proposed association score in (21) is justified.
@@
Figure 7: Showing that the association score in (21) allows track identities to be maintained while targets leave the ground plane. The list displayed in purple represents the model probabilities of the image plane and ground plane filters, respectively, for each track. As targets leave the ground plane, the image plane filter starts to dominate the total model probability.
@@ 6 Results
This section presents the results of IMM-JHSE evaluated on the DanceTrack, MOT17, MOT20 and KITTI test datasets. The results for the DanceTrack test set are shown in Table 4, where IMM-JHSE significantly outperforms related methods in terms of HOTA, IDF1 and AssA, using the same detections. Specifically, compared to the runner-up (UCMCTrack+), HOTA increases by $2.64$, IDF1 by $6.72$ and Assa by $4.11$. Following the ablation study presented in Table 1, this is attributed to IMM-JHSE’s dynamic noise estimation, ground plane IMM filter, which accounts for two camera motion models, and its specific use of BIoU during association. Note that MOTA is the only metric that has not improved over previous methods. It is possible that this is due to the fact that tentative tracks are removed from the result file if they are never confirmed. Thus increasing the number of false negative detections and reducing the MOTA. It is also possible that partial detections are associated with IMM-JHSE, which do not have corresponding ground truth detections because the object is, for example, mostly occluded. This would increase the number of false positive detections, which could also cause a decrease in MOTA. The last row of Table 4 shows the results of FusionTrack [57] on the DanceTrack test set. FusionTrack performs joint detection and tracking with an attention-based decoder and also uses appearance features generated from an encoder-like component. Despite this, IMM-JHSE outperforms FusionTrack in terms of HOTA.
@@
Table 4: Results on the DanceTrack test set with detections obtained from [9], except for FusionTrack [57], which performs tracking and detection jointly with an attention-based architecture.
IMM-JHSE achieves the results reported in Table 5 on the MOT17 test dataset. While IMM-JHSE does not achieve the best result for any metric, it places itself firmly in between UCMCTrack and UCMCTrack+ for all metrics – outperforming all other methods (besides UCMCTrack+) in HOTA, IDF1 and AssA. As will be seen in Tables 6 and 7, IMM-JHSE seems to perform poorly with tracks which take up relatively little of the image plane area, particularly if a multitude of such tracks are in close proximity to one another. In other words, the merits of IMM-JHSE are more evident when tracking relatively large (in terms of bounding box area) objects that move irregularly or with high acceleration or velocity in the image plane rather than in highly cluttered situations. This may be due to the second association stage, which is designed to account for ambiguous detections that may diverge from that expected by the predicted ground plane state. Yet, if multiple tracks are close to one another and all have ambiguous detections, it is conceivable that they may be confused with one another since they are likely to have high BIoU scores with detections that belong to other tracks.
@@
Table 5: Results on the MOT17 test set with detections obtained from [9].
Since MOT20 deals with even more cluttered environments than MOT17, this reasoning seems all but confirmed in Table 6, where IMM-JHSE achieves the worst results for every metric except AssA on the MOT20 test dataset. Another possibility is that the optimised parameters overfit the validation sets, reducing performance on the test sets. Indeed, the pattern search algorithm seems prone to getting stuck in local minima based on close observation during optimisation iterations.
@@
For the KITTI test dataset, IMM-JHSE once again shows the poorest results for the pedestrian class in all metrics except AssA in Table 7. However, it outperforms the HOTA and AssA scores of other methods in the car class – with HOTA increasing by $2.11$ over UCMCTrack and AssA by $5.09$.
Method Car Pedestrian
OC-SORT [11] 76.5 90.3 76.4 54.7 65.1 59.1
UCMCTrack [14] 77.1 90.4 77.2 55.2 67.4 58.0
UCMCTrack+ [14] 74.2 90.2 71.7 54.3 67.2 56.3
IMM-JHSE (Ours) 79.21 89.8 82.29 54.07 64.95 56.89
@@ 7 Conclusion
This work introduced IMM-JHSE, a novel MOT algorithm. The method makes use of an IMM ground plane filter, which combines static and dynamic camera motion models. Crucially, object motion is decoupled from these motion models. Furthermore, image plane information is incorporated through a simple bounding box motion filter and the BIoU score. IMM-JHSE dynamically switches between and mixes BIoU and ground plane-based Mahalanobis distance in an IMM-like fashion to perform association. Finally, dynamic process and measurement noise estimation are used. All of these components contribute to the overall performance of the method.
IMM-JHSE improves significantly upon similar methods on the DanceTrack and KITTI-car datasets while offering competitive performance on the MOT17, MOT20 and KITTI-pedestrian datasets. Future work may improve its performance on data with high track densities.

### v2

@@ 5.1 Parameter selection
The largest assumption of IMM-JHSE is that an initial homography estimate is available. For all experiments and test set results, the camera calibration matrices provided in [32] are used – except for the KITTI dataset, which provides its own. Furthermore, the estimates of $\mathbf{\hat{A}}_{t}$ (2) provided by [32] are used for the MOT17 and MOT20 datasets, but for the DanceTrack and KITTI datasets $\mathbf{\hat{A}}_{t}$ is estimated as mentioned in Section 3.
@@
For IMM-JHSE, the following hyperparameters need to be specified: $\sigma_{x},\sigma_{y},\alpha_{1},\alpha_{2},\alpha_{3},\Omega,b,d_{\text{high}},d_{\text{low}}$, as well as the transition probabilities illustrated in Figure 5. For the transition probabilities, it is only necessary to specify the self-transition probabilities $p_{s,s},p_{d,d},p_{W,W}$ and $p_{I,I}$, since the other transition probabilities can be derived from them.
5.1.1 Filter initialisation
In addition to the parameters above, the covariance matrices with which each filter is initialised must be specified. The initial covariance matrix associated with the ground plane coordinates $\mathbf{x}^{W}_{t}$ is obtained similar to [32]:
$\mathbf{J^{IW}_{t}}\mathbf{R}_{0}\left(\mathbf{J^{IW}_{t}}\right)^{T},$ (32)
where $\mathbf{J^{IW}_{t}}$ is obtained from (19) and $\mathbf{R}_{0}$ from (16). The variance in each velocity component is set to $v$, which is determined with the pattern search algorithm as described subsequently. Finally, the initial covariance associated with the homography elements is set to $\mathbf{0}$. Thus, a perfect homography estimate is assumed initially, but this covariance is dynamically adjusted by (26).
@@ 5.1.1 Filter initialisation
This paper uses the pattern search algorithm [83] to find optimum values for the hyperparameters. For each experiment described subsequently, the initial values are set as follows: $\sigma_{x}=5,\sigma_{y}=5,\alpha_{1}=\alpha_{2}=\alpha_{3}=0.5,\Omega=30,b=0,d_{\text{high}}=0.6,d_{\text{low}}=0.5,p_{s,s}=p_{d,d}=p_{W,W}=p_{I,I}=0.9,v=0.5$, and the number of pattern search iterations is restricted to 200. The objective function selected for optimisation is HOTA for the MOT17, MOT20 and KITTI datasets and association accuracy (AssA) for the DanceTrack dataset. Other evaluation metrics include IDF1 and multiple object tracking accuracy (MOTA). The reader is referred to [84] for details regarding these metrics. In all cases, evaluation is performed with the TrackEval library [85]. For the DanceTrack, KITTI and MOT20 datasets, the average of the objective functions of each video in the entire dataset is optimised, while for the MOT17 dataset, the objective function is optimised per video.
The parameters obtained for IMM-JHSE with the pattern search algorithm are given in Table 9.
5.2 Ablation study on the DanceTrack dataset
@@
Table 1 reports the results of an ablation study on the DanceTrack validation dataset. The second column refers to using the dynamic measurement noise estimation method described in Section 4.3. When dynamic measurement noise is not used (as indicated with X), (16) is used to provide an estimate of the measurement noise at each time step. When the dynamic estimation method is not used, HOTA decreases from $63.81$ to $60.80$. Thus, the dynamic noise estimation method contributes significantly to the success of IMM-JHSE.
The third column refers to the static camera motion model (11). IMM-JHSE is restricted to the dynamic camera motion model of (10) without the static camera motion model. In this case, HOTA decreases to $62.04$ – showing that the static camera motion model (11) is beneficial.
The fourth column refers to using the dynamic camera motion model (10). Without this model, HOTA decreases to $61.19$. Therefore, the dynamic camera motion model plays a more significant role than the static camera motion model.
The fifth column refers to using the image plane filter and the BIoU. Without these, all association scores are calculated by $P(D)d_{\text{conf}}$, and HOTA decreases to $63.03$. Thus, using bounding box information with the ground plane state vector proves beneficial.
The best results are obtained in the first row when all of the components of IMM-JHSE are in use.
@@
5.3 Investigating association score mixture methods
In Table 2, the way bounding box information is considered is varied. For “IMM-like”, the association score for stages 2 and 3 remains as in (30). For “Multiplicative”, all stages use the association score in (27). The “IMM-like” association scores achieve the best performance across all metrics for MOT17 and DanceTrack. Furthermore, the benefit of using the “IMM-like” association score is illustrated in Figure 7, where track identities are maintained while the performers jump, and the ground plane position estimates cannot be relied upon. Thus, the proposed association score in (30) is justified.
@@ 5.3 Investigating association score mixture methods
Figure 7: Showing that the association score in (30) allows track identities to be maintained while targets leave the ground plane. The list displayed in purple represents the model probabilities of the image plane and ground plane filters, respectively, for each track. The image plane filter dominates the total model likelihood as targets leave the ground plane.
5.4 Investigation of the chosen objective function
@@ 6 Results
This section presents the results of IMM-JHSE evaluated on the DanceTrack, MOT17, MOT20 and KITTI test datasets.
6.1 Results on DanceTrack
The results for the DanceTrack test set are shown in Table 4, where IMM-JHSE significantly outperforms related methods (ByteTrack, C-BIoU, OC-SORT and UCMCTrack) in terms of HOTA, IDF1 and AssA, using the same detections. Specifically, compared to the runner-up (UCMCTrack+), HOTA increases by $2.64$, IDF1 by $6.72$ and AssA by $4.11$. Following the ablation study presented in Table 1, this is attributed to IMM-JHSE’s dynamic noise estimation, ground plane IMM filter, which accounts for two camera motion models, and its specific use of BIoU during association. Note that MOTA is the only metric that has not improved over previous methods. This may be because tentative tracks are removed from the result file if they are never confirmed. Thus increasing the number of false negative detections and reducing the MOTA. It is also possible that partial detections are associated with IMM-JHSE, which do not have corresponding ground truth detections because the object is, for example, mostly occluded. This would increase the number of false positive detections, which could also cause a decrease in MOTA.
IMM-JHSE performs remarkably well compared to attention-based methods. Despite FusionTrack’s ability to perform joint detection, tracking, and exploit appearance features, IMM-JHSE outperforms it in terms of HOTA. IMM-JHSE performs similarly to the computationally expensive MeMOTR with Deformable DETR, even outperforming it with IDF1 and association accuracy. IMM-JHSE also outperforms MeMOTR using DAB-Deformable DETR in IDF1. Considering the vast difference between the methods of IMM-JHSE and MOTIP with Deformable DETR, it is remarkable that IMM-JHSE obtains similar performance in most metrics. However, the advantage of the end-to-end approach proves to be great with the combination of MOTIP and DAB-Deformable DETR, which significantly outperforms IMM-JHSE.
@@ 6.1 Results on DanceTrack
MeMOTR [45] 66.7 70.6 94 53
MeMOTR [45] 68.5 71.2 89.9 58.4
MOTIP [46] 67.5 72.2 90.3 57.6
MOTIP [46] 70 75.1 91 60.8
Deformable DETR [86]
DAB-Deformable DETR [87]
Table 4: Results on the DanceTrack test set. For the tracking-by-detection methods, the detections are obtained from [9].
6.2 Results on the KITTI dataset
For the KITTI test dataset, IMM-JHSE significantly outperforms the HOTA and AssA scores of other methods (UCMCTrack and OC-SORT) in the car class – with HOTA increasing by $2.11$ over UCMCTrack and AssA by $5.09$ in Table 5.
At the time of writing, IMM-JHSE is placed twenty-fourth on the KITTI-car leaderboard [^1] https://www.cvlibs.net/datasets/kitti/eval_tracking_detail.php?result=c3e8939f18cba7b23a2e05a92d4df9e994eaaf91. In addition, all methods that are placed above it are 3D MOT methods that make use of regular 3D measurements, except for RAM [88], which uses a self-supervised encoder that optimises a Markov walk along a space-time graph. Table 6 compares IMM-JHSE and some of these methods. Interestingly, while RAM outperforms IMM-JHSE in the car class, IMM-JHSE outperforms RAM in the pedestrian class. While there is a large difference (3.49 HOTA) between IMM-JHSE and the top-ranked method (BiTrack [19]), it should also be noted that IMM-JHSE uses publicly available detections and the corresponding difference in detection accuracy is 3.64%. Furthermore, BiTrack is an offline method.
However, IMM-JHSE shows the poorest results for the pedestrian class in all metrics except AssA in Table 5. As will be seen in Table 7 and Table 8 for the MOT17 and MOT20 datasets, IMM-JHSE seems to perform poorly with tracks which take up relatively little of the image plane area, particularly if a multitude of such tracks are close to one another. In other words, the merits of IMM-JHSE are more evident when tracking relatively large (in terms of bounding box area) objects that move irregularly or with high acceleration or velocity in the image plane rather than in highly cluttered situations. This may be due to the second association stage, which is designed to account for ambiguous detections that may diverge from that expected by the predicted ground plane state. Yet, suppose multiple tracks are close to one another, and all have ambiguous detections. In that case, it is conceivable that they may be confused with one another since they are likely to have high BIoU scores with detections that belong to other tracks.
Another possibility is that the optimised parameters overfit the validation set on the car class, reducing performance on the pedestrian test set. Indeed, the pattern search algorithm seems prone to getting stuck in local minima based on close observation during optimisation iterations.
Despite the relatively poor performance of IMM-JHSE with the pedestrian class, it is placed third on the KITTI-pedestrian leaderboard at the time of writing, where OC-SORT is the only published method that outperforms it.
Method Car Pedestrian
OC-SORT [11] 76.5 90.3 76.4 54.7 65.1 59.1
UCMCTrack [32] 77.1 90.4 77.2 55.2 67.4 58.0
UCMCTrack+ [32] 74.2 90.2 71.7 54.3 67.2 56.3
IMM-JHSE (Ours) 79.21 89.8 82.29 54.07 64.95 56.89
Method Car Pedestrian
AB3DMOT
+PointRCNN [21] 69.99 83.61 69.33 37.81 38.13 44.33
Mono 3D KF [20] 75.47 88.48 77.63 42.87 45.44 46.31
OC-SORT [11] 76.54 90.28 76.39 54.69 65.14 59.08
PermaTrack [89] 78.03 91.33 78.41 48.63 65.98 45.61
FusionTrack
+pointgnn [44] 78.50 89.57 81.89
IMM-JHSE (Ours) 79.21 89.8 82.29 54.07 64.95 56.89
RAM [88] 79.53 91.61 80.94 52.71 68.4 52.19
MCTrack (online) [26] 80.78 89.82 84.30
CasTrack [17] 81.00 91.91 84.22
MCTrack (offline) [26] 82.56 91.62 86.64
BiTrack (offline) [19] 82.7 91.55 86.17
Table 6: A comparison of IMM-JHSE to other methods on the KITTI leaderboard.
6.3 Results on the MOT datasets
IMM-JHSE achieves the results reported in Table 7 on the MOT17 test dataset [^2] https://motchallenge.net/method/MOT=8468&chl=10. While IMM-JHSE does not achieve the best result for any metric, it places itself firmly in between UCMCTrack and UCMCTrack+ for all metrics—outperforming all other methods besides UCMCTrack+—including ByteTrack, C-BIoU and OC-SORT—in HOTA, IDF1 and AssA.
IMM-JHSE also significantly outperforms tracking-by-attention methods on the MOT17 dataset. While the performance of IMM-JHSE may be disappointing, it is clear that the high track densities in the MOT datasets also present a challenge to tracking-by-attention methods.
@@ 6.3 Results on the MOT datasets
Tracking-by-attention
MeMOTR [45] 58.8 71.5 72.8 58.4
MOTIP [46] 59.2 71.2 75.5 56.9
Deformable DETR [86]
Table 7: Results on the MOT17 test set. For the tracking-by-detection methods, the detections are obtained from [9].
Since MOT20 deals with even more cluttered environments than MOT17, it is unsurprising that IMM-JHSE achieves the worst results for every metric except AssA on the MOT20 test dataset in Table 8.
@@ 7 Conclusion
This work introduced IMM-JHSE, a novel MOT algorithm. The method uses an IMM ground plane filter, which combines static and dynamic camera motion models. Crucially, object motion is decoupled from these motion models. Furthermore, image plane information is incorporated through a simple bounding box motion filter and the BIoU score. IMM-JHSE dynamically switches between and mixes BIoU and ground plane-based Mahalanobis distance in an IMM-like fashion to perform association. Finally, dynamic process and measurement noise estimation are used. All of these components contribute to the overall performance of the method.
IMM-JHSE performs remarkably well compared to fully 3D MOT methods despite only using a single initial homography estimate, especially considering some 3D MOT methods on the KITTI leaderboard are offline and/or use their own detections.
Furthermore, IMM-JHSE performs remarkably similarly to attention-based methods on the DanceTrack dataset and outperforms them on the MOT17 dataset. Not only does IMM-JHSE bridge the gap between 3D and 2D MOT methods, but also between tracking-by-detection and tracking-by-attention methods.
IMM-JHSE improves significantly upon related methods with the same publicly available detections, such as UCMCTrack, C-BIoU, OC-SORT and ByteTrack, on the DanceTrack and KITTI-car datasets while offering competitive performance on the MOT17, MOT20 and KITTI-pedestrian datasets.
However, a weak point of IMM-JHSE is that it performs poorly when tracks are very dense, especially in the MOT20 dataset. This is attributed to the leniency afforded by the multitude of motion models in use. Future work may, therefore, improve its performance on data with high track densities.
@@ Acknowledgements
Appendix A IMM filter theory
IMM-JHSE runs two ground plane EKFs in parallel, each with the state vector definition in Equation 13. One uses the static homography dynamics model in Equation 11, while the other uses the dynamic one in Equation 10. The IMM methodology allows the proposed method to switch between and mix state estimates from both filters, proportional to their likelihood [80]. This way, IMM-JHSE is robust to scenes with both static and dynamic camera movement. This section briefly describes the operation of IMM filters according to [77, 79, 78, 80], which is implemented with the FilterPy library [90].
Consider a system with $r$ different models. The following state-space equations describe each model $i$ (assumed linear for simplicity):
$\begin{split}\mathbf{x}_{t}^{i}&=\mathbf{A}_{t}^{i}\mathbf{x}_{t-1}^{i}+\mathbf{B}_{t}^{i}\mathbf{u}_{t}^{i}+\mathbf{w}_{t}^{i},\\ \mathbf{y}_{t}^{i}&=\mathbf{G}_{t}^{i}\mathbf{x}_{t}^{i}+\mathbf{v}_{t}^{i},\end{split}$
where $\mathbf{x}_{t}^{i}$ is the state vector for model $i$ at time $t$, $\mathbf{A}_{t}^{i}$, $\mathbf{B}_{t}^{i}$, and $\mathbf{G}_{t}^{i}$ are the state transition, control input, and observation matrices, respectively. $\mathbf{u}_{t}^{i}$ is the control input and $\mathbf{w}_{t}^{i}$ and $\mathbf{v}_{t}^{i}$ are process and measurement noise, respectively.
The probability of each model $i$ being the correct model at time $t$ is denoted by $\mu_{t}^{i}$. These probabilities are updated at each time step based on the likelihood of the observations given each model.
The mixing probabilities $\mu_{t|t-1}^{i|j}$ represent the probability of transitioning from model $j$ at time $t-1$ to model $i$ at time $t$. These are computed using the Markov transition probabilities $p_{ji}$:
$\mu_{t|t-1}^{i|j}=\frac{p_{ji}\mu_{t-1}^{j}}{\sum_{l=1}^{r}p_{li}\mu_{t-1}^{l}}.$
The mixed initial conditions for each filter are computed as follows:
$\begin{split}\hat{\mathbf{x}}_{t-1|t-1}^{i}&=\sum_{j=1}^{r}\mu_{t|t-1}^{i|j}\hat{\mathbf{x}}_{t-1|t-1}^{j},\\ \mathbf{P}_{t-1|t-1}^{i}&=\sum_{j=1}^{r}\mu_{t|t-1}^{i|j}\left(\mathbf{P}_{t-1|t-1}^{j}+(\hat{\mathbf{x}}_{t-1|t-1}^{j}-\hat{\mathbf{x}}_{t-1|t-1}^{i})(\hat{\mathbf{x}}_{t-1|t-1}^{j}-\hat{\mathbf{x}}_{t-1|t-1}^{i})^{T}\right).\end{split}$
Each model-specific Kalman filter is updated using the mixed initial conditions:
$\begin{split}\hat{\mathbf{x}}_{t|t-1}^{i}&=\mathbf{A}_{t}^{i}\hat{\mathbf{x}}_{t-1|t-1}^{i}+\mathbf{B}_{t}^{i}\mathbf{u}_{t}^{i},\\ \mathbf{P}_{t|t-1}^{i}&=\mathbf{A}_{t}^{i}\mathbf{P}_{t-1|t-1}^{i}(\mathbf{A}_{t}^{i})^{T}+\mathbf{Q}_{t}^{i},\\ \mathbf{K}_{t}^{i}&=\mathbf{P}_{t|t-1}^{i}(\mathbf{G}_{t}^{i})^{T}(\mathbf{G}_{t}^{i}\mathbf{P}_{t|t-1}^{i}(\mathbf{G}_{t}^{i})^{T}+\mathbf{R}_{t}^{i})^{-1},\\ \hat{\mathbf{x}}_{t|t}^{i}&=\hat{\mathbf{x}}_{t|t-1}^{i}+\mathbf{K}_{t}^{i}(\mathbf{y}_{t}-\mathbf{G}_{t}^{i}\hat{\mathbf{x}}_{t|t-1}^{i}),\\ \mathbf{P}_{t|t}^{i}&=(\mathbf{I}-\mathbf{K}_{t}^{i}\mathbf{G}_{t}^{i})\mathbf{P}_{t|t-1}^{i}.\end{split}$
The model probabilities are updated based on the likelihood of the observations:
$\mu_{t}^{i}=\frac{\Lambda_{t}^{i}\sum_{j=1}^{r}p_{ji}\mu_{t-1}^{j}}{\sum_{l=1}^{r}\Lambda_{t}^{l}\sum_{j=1}^{r}p_{jl}\mu_{t-1}^{j}},$
where $\Lambda_{t}^{i}$ is the likelihood of the observation given model $i$.
The overall state estimate and covariance are computed as a weighted sum of the individual model estimates:
$\begin{split}\hat{\mathbf{x}}_{t}&=\sum_{i=1}^{r}\mu_{t}^{i}\hat{\mathbf{x}}_{t|t}^{i},\\ \mathbf{P}_{t}&=\sum_{i=1}^{r}\mu_{t}^{i}\left(\mathbf{P}_{t|t}^{i}+(\hat{\mathbf{x}}_{t|t}^{i}-\hat{\mathbf{x}}_{t})(\hat{\mathbf{x}}_{t|t}^{i}-\hat{\mathbf{x}}_{t})^{T}\right).\end{split}$
Appendix B Deriving the Jacobian matrix
Let the vector $\mathbf{b}$ represent the unnormalised projection of a ground plane coordinate, i.e.
[…]
where $x^{W}_{t}$ and $y^{W}_{t}$ represent the x- and y-coordinate of the ground plane coordinate under consideration, respectively. Expanding the above with
$\mathbf{H}_{t}=\begin{bmatrix}h_{1}&h_{2}&h_{3}\\ h_{4}&h_{5}&h_{6}\\ h_{7}&h_{8}&h_{9}\\ \end{bmatrix}:$
$\begin{bmatrix}b_{1}\\ b_{2}\\ b_{3}\end{bmatrix}=\begin{bmatrix}h_{1}x^{W}_{t}+h_{2}y^{W}_{t}+h_{3}\\ h_{4}x^{W}_{t}+h_{5}y^{W}_{t}+h_{6}\\ h_{7}x^{W}_{t}+h_{8}y^{W}_{t}+h_{9}\end{bmatrix}.$
The Jacobian of $\mathbf{b}$ with respect to the world coordinates is:
$\begin{bmatrix}\frac{\partial b_{1}}{\partial x^{W}_{t}}&\frac{\partial b_{1}}{\partial y^{W}_{t}}\\ \frac{\partial b_{2}}{\partial x^{W}_{t}}&\frac{\partial b_{2}}{\partial y^{W}_{t}}\\ \frac{\partial b_{3}}{\partial x^{W}_{t}}&\frac{\partial b_{3}}{\partial y^{W}_{t}}\end{bmatrix}=\begin{bmatrix}h_{1}&h_{2}\\ h_{4}&h_{5}\\ h_{7}&h_{8}\end{bmatrix}.$
Note that the image coordinates are obtained by $\begin{bmatrix}x^{I}_{t}&y^{I}_{t}\end{bmatrix}^{\top}=\begin{bmatrix}b_{1}/b_{3}&b_{2}/b_{3}\end{bmatrix}^{\top}$. Therefore, the partial derivative of $x^{I}_{t}$ with respect to $x^{W}_{t}$ is:
$\frac{\partial x^{I}_{t}}{\partial x^{W}_{t}}$ $=\frac{\partial b_{1}b_{3}^{-1}}{\partial x^{W}_{t}}$
$=\frac{h_{1}b_{3}-h_{7}b_{1}}{b_{3}^{2}}$
$=\gamma(h_{1}-h_{7}x^{I}_{t}),$
where $\gamma=b_{3}^{-1}$. Similarly,
$\frac{\partial x^{I}_{t}}{\partial y^{W}_{t}}$ $=\gamma(h_{2}-h_{8}x^{I}_{t}),$
$\frac{\partial y^{I}_{t}}{\partial x^{W}_{t}}$ $=\gamma(h_{4}-h_{7}y^{I}_{t}),$
$\frac{\partial y^{I}_{t}}{\partial y^{W}_{t}}$ $=\gamma(h_{5}-h_{8}y^{I}_{t}).$
Thus, the Jacobian matrix of the x- and y- image coordinates with respect to the x- and y- world coordinates is
$\mathbf{J}^{IW}_{t}=\begin{bmatrix}\frac{\partial x^{I}_{t}}{\partial x^{W}_{t}}&\frac{\partial x^{I}_{t}}{\partial y^{W}_{t}}\\ \frac{\partial y^{I}_{t}}{\partial x^{W}_{t}}&\frac{\partial y^{I}_{t}}{\partial y^{W}_{t}}\end{bmatrix}=\gamma\begin{bmatrix}(h_{1}-h_{7}x^{I}_{t})&(h_{2}-h_{8}x^{I}_{t})\\ (h_{4}-h_{7}y^{I}_{t})&(h_{5}-h_{8}y^{I}_{t})\end{bmatrix}.$
Similarly, the Jacobian matrix of the x- and y- image coordinates with respect to the homography elements is
$\mathbf{J}^{IH}_{t}$ $=\begin{bmatrix}\frac{\partial x^{I}_{t}}{\partial h_{1}}&\frac{\partial x^{I}_{t}}{\partial h_{4}}&\frac{\partial x^{I}_{t}}{\partial h_{7}}&\frac{\partial x^{I}_{t}}{\partial h_{2}}&\frac{\partial x^{I}_{t}}{\partial h_{5}}&\frac{\partial x^{I}_{t}}{\partial h_{8}}&\frac{\partial x^{I}_{t}}{\partial h_{3}}&\frac{\partial x^{I}_{t}}{\partial h_{6}}&\frac{\partial x^{I}_{t}}{\partial h_{9}}\\ \frac{\partial y^{I}_{t}}{\partial h_{1}}&\frac{\partial y^{I}_{t}}{\partial h_{4}}&\frac{\partial y^{I}_{t}}{\partial h_{7}}&\frac{\partial y^{I}_{t}}{\partial h_{2}}&\frac{\partial y^{I}_{t}}{\partial h_{5}}&\frac{\partial y^{I}_{t}}{\partial h_{8}}&\frac{\partial y^{I}_{t}}{\partial h_{3}}&\frac{\partial y^{I}_{t}}{\partial h_{6}}&\frac{\partial y^{I}_{t}}{\partial h_{9}}\\ \end{bmatrix}$
$=\gamma\begin{bmatrix}x^{W}_{t}&0&-x^{I}_{t}x^{W}_{t}&y^{W}_{t}&0&-x^{I}_{t}y^{W}_{t}&1&0&-x^{I}_{t}\\ 0&x^{W}_{t}&-y^{I}_{t}x^{W}_{t}&0&y^{W}_{t}&-y^{I}_{t}y^{W}_{t}&0&1&-y^{I}_{t}\end{bmatrix}.$
Finally, the combined Jacobian matrix of the image coordinates with respect to the homography and world coordinates is given by
$\mathbf{J}_{t}=\begin{bmatrix}\mathbf{J}^{IW}_{t}&\mathbf{0}\\ \mathbf{0}&\mathbf{J}^{IH}_{t}\end{bmatrix}.$
Appendix C Parameter results
Table 9 shows the parameters obtained with the pattern search algorithm for the various datasets considered. For the DanceTrack dataset, these are the parameters obtained by optimising the overall AssA. For the KITTI, MOT17 and MOT20 datasets, these are the parameters obtained by optimising the overall HOTA.
Parameter DanceTrack MOT17 MOT20 KITTI
$\sigma_{x}$ 5 5 0.001 14.84
$\sigma_{y}$ 18.75 3.6 5.0 16.25
$\alpha_{1}$ 0.27 0.91 0.91 0.83
$\alpha_{2}$ 0.99 0.46 0.5 0.5
$\alpha_{3}$ 0.82 0.99 0.91 0.91
$\Omega$ 62 30 62 100
$b$ 0 0 0 0
$d_{\text{high}}$ 0.9 0.6 0.75 0.29
$d_{\text{low}}$ 0.1 0.26 0.5 0.5
$p_{s,s}$ 0.16 0.99 0.9 0.72
$p_{d,d}$ 0.53 0.9 0.99 0.72
$p_{W,W}$ 0.9 0.9 0.72 0.9
$p_{I,I}$ 0.01 0.9 0.53 0.9
$v$ 0.56 $1\times 10^{-7}$ 0.005 0.009
Table 9: Parameters of IMM-JHSE for different datasets.